OpenStudy (anonymous):

Write the given equation in standard form and determine the center and radius. x2 + 4x + y2 – 10y – 35 = 0

OpenStudy (anonymous):

you'll need to complete the square to get it into the standard form: (x-h)^2 + (y-k)^2 = r^2 to find the center (h,k) and radius r. complete the square on: x^2 + 4x + y^2 – 10y – 35 = 0 (x^2 + 4x ) + (y^2 – 10y ) =35 what do you need to add to each of those parenthesis to make each a perfect square trinomial?

OpenStudy (anonymous):

idk what?

OpenStudy (anonymous):

take half of the linear coefficient and square it. that is what you need to add to each to make each parenthesis into a perfect square trinomial. what's the coefficient of the linear term (x), in the first parenthesis?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

no. the coefficient of the linear term is 4... take half of that, then square it: 4/2=2 ---> 2^2=4 this is what you need to add to the first parenthesis. (x^2 + 4x + 4) + (y^2 – 10y ) =35 can you find out what you need to add to make that second parenthesis a perfect square trinomial....

OpenStudy (anonymous):

4?

OpenStudy (anonymous):

(x^2 + 4x + 4) + (y^2 – 10y ) =35 + 4 (x^2 + 4x + 4) + (y^2 – 10y + 25 ) =35 + 4 + 25 (x+2)^2 + (y-5)^2 = 64 ok.. i completed it for you. you need to determine the center and radius from here. that last equation contains everything you need to know about the center and radius...:)